### v1

# Mellin‑Space Prony Representability of Linear Viscoelastic Models

Abstract
Linear viscoelasticity is universally described by relaxation functions, which are continuous in relaxation times. Yet experimentally, measurements are necessarily discrete and band-limited, creating a fundamental gap between the continuous mathematical description and empirical observations. While Laplace-domain rational moduli admit finite Prony representations, using Mellin space analysis requires a deeper structural criterion. This work proves that finite Prony series representation holds if and only if the arithmetic pole lattices of Gamma factors in the Mellin transform of the complex modulus align exactly with those of a trial kernel, with residues satisfying decoupled first-order recurrences along aligned sublattices. This maximal criterion classifies classical models (Maxwell, standard linear solid) as finitely representable and fractional models (power-law, Cole–Cole, Zener, Gaussian) as transcendentally representable via infinite Prony ladders, yielding a complete analytical pole-lattice taxonomy of viscoelastic material models.
keywords:
Mellin transform , linear viscoelasticity , Prony series , relaxation spectra , fractional calculus
PACS:
83.60.Bc , 02.30.Uu , 46.35.+z , 02.30.Gp , 02.30.Hq
2020 MSC:
44A20 , 33E12 , 33B15 , 30E20 , 74D05
journal: Chaos Solitions and Fractals
[…]
organization=Laboratory of Neurotechnology, PAML-LN, Inistitute of Information and Communication Technologies, Bulgarian Academy of Science,addressline=2, acad. G.Bontchev Str, bl. 25A, city=Sofia, postcode=1113, state=, country=Bulgaria
1 Introduction
A material is linear viscoelastic if stress is a linear time-dependent functional of strain. Linear viscoelasticity is universally described by relaxation functions, which are continuous distributions of relaxation times. Yet experimentally, measurements are necessarily discrete and band-limited, creating a fundamental gap between the continuous mathematical description and empirical observations.
With the caveat that “all models are wrong but some are useful” [1] – a stance universally adopted in statistics and data science —this works explores the utility of Prony series in describing linear viscoelasticty from complex-analytical perspective. The Prony series provides an approximate representation of any viscoelastic model through finite sums of exponential terms in its time-dependent relaxation moduli. A comprehensive recent review of Prony-series applications in viscoelasticity is given in [2].
From a dynamical-systems perspective, finding a linear viscoelastic representation is equivalent to constructing a mechanical network—a finite arrangement of springs and dashpots—whose input-output behavior reproduces the prescribed complex=valued dynamic modulus $G^{*}(\omega)$. Such networks generate relaxation spectra that are finite sums of Dirac masses (finite Prony series) with rational transfer functions in the frequency (Fourier/Laplace) variable. The inverse question—whether a given modulus admits a finite network representation—therefore translates into a classical realization problem: does the system possess a finite-dimensional state-space description, or equivalently, does its transfer function exhibit a pole structure consisting of a finite set of simple poles on the negative real axis? The Mellin-transform framework developed here recasts this condition as an algebraic constraint on the pole lattices of the transformed kernel, thereby linking the geometry of the complex-plane singularities directly to the existence of finite spring–dashpot analogous networks.
Classical spring-dashpot networks (Maxwell, Kelvin-Voigt, Standard Linear Solid, and their finite generalizations) yield linear ordinary differential equations for the constitutive relations. Furthermore, their complex moduli $G^{*}(\omega)$ are rational functions of the imaginary spectral variable $i\omega$ in the Fourier/Laplace domains, admitting exact Prony representations with finitely many exponential terms. That is, in the frequency domain their spectrum contains a finite number of poles. Fractional constitutive laws, by contrast, involve fractional memory-dependent, non-local (e.g. containing combinations of differential and integral operators, i.e. ”differintergrals”) constitutive laws [3]. Many soft materials exhibit power-law viscoelastic responses across typical observation bandwidths, modeled by continuous distributions of relaxation times that reflect their fractal internal structure [4]). Structurally, power-law models, Cole-Cole, and Havriliak-Negami kernels generate non-rational dynamic moduli $G^{*}(\omega)$ with branch-cuts, implying infinite relaxation times distributions.
Recent works surveys Prony-series approximations of fractional materials’ data  [2, 5]. Yet, these approximations remain ad hoc constructions for a given model. To understand the dichotomy between finite Prony models and non-local fractional models one needs to resolve the inverse problem: given a modeled $G^{*}(\omega)$, does it admit a faithful finite Prony representation? This question is numerically ill-posed: recovering $\{g_{k},\tau_{k}\}$ from frequency data suffers Hadamard instability, where noise amplifies into spurious oscillations, obscuring the discrete-vs-continuous distinction [6, 7]. Therefore, to establish the answer one needs to employ analytical tools.
We resolve so-stated dichotomy through Mellin-space analysis. In the Mellin complex plane fractional kernels possess kernels (i.e. Mellin symbols) containing non-degenerate Gamma factors, whose poles form infinite arithmetic progressions. We prove that exact finite Prony representability holds if and only if these Gamma-factor pole lattices align exactly with trial state’s spectrum poles, with residues satisfying decoupled first-order recurrence relations along sublattices. This maximal criterion classifies classical models (Maxwell, SLS) as finitely representable and fractional models (power-law, Cole-Cole, Gaussian) as transcendentally representable via infinite Prony ladders.
The paper is organized as follows. General notation and conventions are stated in Section 2. Section 3 introduces the Mellin‑space formulation of linear viscoelasticity. Section 4 states and proves the main criterion—a necessary and sufficient condition for finite representability within the maximal Prony class $\mathcal{P}$. Section 6 classifies common viscoelastic models according to this criterion. Section 7 provides a practical computational test for finite representability. Technical details are relegated to the Appendix to maintain the flow of presentation.
2 Notation and Conventions
2.1 Mellin transforms
For a function $f(t)$ integrable on the positive half-line $(0,\infty)$ whose growth at about 0 and infinity is at most power-law, the Mellin transform is defined by the complex-valued integral
$\mathcal{M}\{f\}(s)\equiv\tilde{f}(s):=\int_{0}^{\infty}t^{\,s-1}f(t)\,dt,$
whenever the integral converges absolutely in some vertical strip $a<\Re s<b$ of the complex plane [8]. For many functions arising in viscoelasticity (e.g., bounded relaxation spectra), such a strip exists.
Extension to generalized functions. The definition extends to distributions via analytic continuation, and to tempered distributions via the theory of generalized functions. In particular:
  • 1.
    For a discrete spectrum coming from finite Prony series $H(\tau)=\sum_{k=1}^{N}g_{k}\tau_{k}\delta(\tau-\tau_{k})$, the Mellin transform is evaluated as the finite sum
    $\tilde{H}(s)=\sum_{k=1}^{N}g_{k}\tau_{k}^{\,s},$
    which is an entire function of exponential type.
  • 2.
    For a pure power-law $f(\tau)\sim\tau^{\beta}$, the Mellin transform is defined via analytic continuation from the strip $(0,1)\to[1,\infty)$
    $\mathcal{M}\{\tau^{\beta}\}(s)=\int_{0}^{1}\tau^{\beta}d\tau=\frac{1}{\beta+s}$
    and corresponds to a simple pole at $s=-\beta$.
All Mellin transforms appearing in this work are either classical (convergent in a strip) or defined via such analytic/distributional extensions, ensuring that the subsequent constitutive identity and pole-lattice analysis are mathematically well-defined.
2.2 The Finite Prony Class $\mathcal{P}$
The canonical Prony model [9],
$G_{P}(t)=G_{\infty}+\sum_{k=1}^{N}g_{k}e^{-t/\tau_{k}},$
represents the stress relaxation as a finite superposition of decreasing exponentials, reflecting the fading-memory principle of Boltzmann’s superposition [10]. Furthermore, Bernstein’s theorem [11] guarantees that completely monotone relaxation moduli $G(t)$ admit continuous spectral representations as non-negative Laplace transforms:
$G(t)=G_{\infty}+\int_{0}^{\infty}e^{-t/\tau}H(\tau)\frac{d\tau}{\tau},\quad H(\tau)\geq 0.$ (1)
This establishes realizability for fading-memory materials but leaves unresolved the structural dichotomy between finite Prony series (discrete spectra) and generic continuous distributions, having combinations of poles and branch-cuts.
We call $\mathcal{P}$ the finite Prony class: the maximal class of relaxation spectra whose Mellin symbols contain finite Gamma-ratios (i.e. of Mellin–Barnes/Fox–$H$ type), for which finite-mode realizability can be decided by the pole-lattice alignment, which will be built in Sec. 4. Formally, $\mathcal{P}$ can be defined as:
Definition 1 (Finite Prony class $\mathcal{P}$).
A function $\tilde{H}(s)$ belongs to the finite Prony class $\mathcal{P}$ if it can be written as
$\tilde{H}(s)=H_{\Gamma}(s)\,Q(s),$
where
$H_{\Gamma}(s):=\frac{\prod_{i=1}^{N}\Gamma(a_{i}s+b_{i})}{\prod_{j=1}^{M}\Gamma(c_{j}s+d_{j})},\qquad a_{i},c_{j}>0;\quad N,M<\infty$
is a finite Gamma-ratio and
$Q(s):=\exp(P(s)),\qquad P(s)\ \text{entire function}.$
The factor $Q(s)$ is included to allow entire, non-vanishing kernels (e.g. with exponential or polynomial growth in $s$); it does not affect pole-lattice locations, which are determined solely by the Gamma ratio $H_{\Gamma}(s)$. The class $\mathcal{P}$ contains as proper subclasses:
  • 1.
    Finite discrete Prony spectra,
  • 2.
    Continuous power‑law spectra (single Gamma factor $\Gamma(s+\beta)$),
  • 3.
    Mittag‑Leffler‑type spectra (ratios of Gamma functions with non‑integer parameters)
and can be considered informally as a minimal extension of the Fox H-functions in the sense of Mathai [12].
[…]
$\Lambda_{i}=\Bigl\{-\frac{b_{i}+k}{a_{i}}\;:\;k=0,1,2,\dots\Bigr\},\qquad\Delta_{i}=\frac{1}{a_{i}}.$
[…]
3 Mellin‑Space Formulation of Linear Viscoelasticity
We adopt the following notation conventions: $t$ denotes physical time, $\tau>0$ represents relaxation timescale, $\omega>0$ is angular frequency, and $s$ is the complex Mellin variable, coupling the relaxation timescale to frequency. The Mellin variable $s$ reveals the hidden logarithmic structure of relaxation spectra, transforming scale-separated Prony series into algebraic pole-lattice geometry.
3.1 Constitutive relation
The non-aging, isothermal stress-strain relation for a linear viscoelastic material is given by the Boltzmann superposition integral:
$\sigma(t)=\int_{-\infty}^{t}G(t-u)\,d\varepsilon(u),$ (2)
where $\sigma(t)$ and $\varepsilon(t)$ are the stress and strain, and $G(t)$ is the relaxation modulus.
In the time domain, the relaxation modulus of a linear viscoelastic solid takes the spectral form (1) as discussed. In the frequency domain, the relation becomes
$\sigma^{*}(\omega)=G^{*}(\omega)\varepsilon^{*}(\omega),$ (3)
where the complex modulus admits the spectral representation
$G^{*}(\omega)=G_{\infty}+\int_{0}^{\infty}\frac{i\omega\tau}{1+i\omega\tau}H(\tau)\frac{d\tau}{\tau}.$ (4)
The spectral kernel is thus identified as
$K(\omega\tau):=\frac{i\omega\tau}{1+i\omega\tau}.$ (5)
As a technical condition we will require that the integral converges absolutely for every $\omega>0$ and that both $G^{*}$ and $H$ possess Mellin transforms in an overlapping vertical strips of the complex plane.
Observe that a finite Prony series corresponds to a discrete spectrum
$H(\tau)=\sum_{k=1}^{N}g_{k}\tau_{k}\delta(\tau-\tau_{k}),$ (6)
producing
$G(t)=G_{\infty}+\sum_{k=1}^{N}g_{k}e^{-t/\tau_{k}}$ (7)
in the time domain.
3.2 Mellin transform of the spectral kernel
The viscoelastic spectral kernel $K(\omega\tau)$ has Mellin transform (w.r.t. $\omega$, fixed $\tau>0$)
$\tilde{K}(s)=\mathcal{M}_{\omega}\{K(\omega\tau)\}(s)=-\pi e^{i\pi(1-s)/2}\csc(\pi s),\qquad 0<\Re s<1,$ (8)
as shown in Appendix Lemma 2.
Remark 1.
The factor $e^{i\pi(1-s)/2}$ is entire (no poles). All singularities come from $\csc(\pi s)$, which has simple poles at $s=n\in\mathbb{Z}$ with residues $(-1)^{n}$. Thus $\tilde{K}(s)$ has simple poles at all integers $s=n$ with residues $e^{i\pi(1-n)/2}(-1)^{n}$.
3.3 Constitutive equation in Mellin space
To avoid dealing with the Dirac $\delta$ distribution let $\Delta G^{*}(\omega)=G^{*}(\omega)-G_{\infty}$, the viscoelastic part of the complex modulus, and
$\tilde{G}(s):=\mathcal{M}\{\Delta G^{*}(\omega)\}(s)$ (9)
be its Mellin transform, and $\tilde{H}(s)=\mathcal{M}\{H(\tau)\}(s)$ be the Mellin transform of the relaxation spectrum. Applying the Mellin transform to (4) and using the convolution theorem for Mellin transforms yields
$\tilde{G}(s)=\tilde{K}(s)\,\tilde{H}(-s),\qquad 0<\Re s<1.$ (10)
Equation (10) is the fundamental Mellin‑space constitutive relation of linear viscoelasticity. It links the unknown spectrum $\tilde{H}(s)$ to the prescribed modulus $\tilde{G}(s)$ through a simple multiplicative kernel that is meromorphic in $s$. Note that since $\tilde{K}(s)$ is odd its argument sign convention corresponds to $\tilde{H}(-s)$. This sign convention preserves the physical duality between the time-scale variable $\tau$ and the frequency $\omega$ in the Mellin representation.
Note also the argument inversion in the Mellin representation $\tilde{H}(-s)$, which determines that $\tilde{H}$ has poles at $\Lambda_{i}^{-}=\{\lambda:-\lambda\in\Lambda_{i}\}$ i.e. the lattices are reflected through the origin.
4 Criteria for Finite Prony Representation
First, we consider the particular case of a degenerate $H$ kernel.
4.1 Degenerate case
Proposition 1 (Characterization of finite Prony spectra).
Let $H(\tau)$ be a positive measure on $(0,\infty)$ with Mellin transform $\tilde{H}(s)$ converging in some vertical strip. Then $H(\tau)$ is a finite Prony series, i.e.,
$H(\tau)=\sum_{k=1}^{N}g_{k}\tau_{k}\delta(\tau-\tau_{k}),\quad g_{k}>0,\ \tau_{k}>0,$
if and only if $\tilde{H}(s)$ is a finite exponential sum of the form
$\tilde{H}(s)=\sum_{k=1}^{N}c_{k}e^{\lambda_{k}s},$
where $c_{k}=g_{k}>0$ and $\lambda_{k}=\ln\tau_{k}\in\mathbb{R}$.
Proof.
($\Rightarrow$) Direct computation gives $\tilde{H}(s)=\sum_{k}g_{k}\tau_{k}^{s}=\sum_{k}g_{k}e^{s\ln\tau_{k}}$.
($\Leftarrow$) Suppose $\tilde{H}(s)=\sum_{k=1}^{N}c_{k}e^{\lambda_{k}s}$ with distinct $\lambda_{k}\in\mathbb{R}$ and $c_{k}\neq 0$. The inverse Mellin transform along a vertical line $\Re s=\sigma$ in the strip of convergence yields
$H(\tau)=\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}\tilde{H}(s)\,ds=\sum_{k=1}^{N}c_{k}\cdot\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}e^{\lambda_{k}s}\,ds.$
The inner integral is the inverse Mellin transform of $e^{\lambda_{k}s}$. Since $e^{\lambda_{k}s}=\mathcal{M}\{\delta(\ln\tau+\lambda_{k})\}(s)$ in the distributional sense, we obtain formally
$\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}e^{\lambda_{k}s}\,ds=\delta(\ln\tau+\lambda_{k})=\tau_{k}\,\delta(\tau-\tau_{k}),$
using the scaling property of the delta function (true in distributional sense)
$\delta(g(\tau))=\sum_{\tau_{i}:\,g(\tau_{i})=0}\frac{\delta(\tau-\tau_{i})}{|g^{\prime}(\tau_{i})|}$
where $\tau_{k}=e^{-\lambda_{k}}$. Hence $H(\tau)=\sum_{k}c_{k}\tau_{k}\delta(\tau-\tau_{k})$, which is a finite Prony series with $g_{k}=c_{k}\tau_{k}$. ∎
4.2 Generic case
To proceed with the analysis we construct a trial state ansatz defined as follows:
Definition 2 (Trial state).
A trial state $G^{*}$ is a modulus which admits a Mellin symbol with factorization:
$\mathcal{M}\{G^{*}(\omega)\}(s)=\tilde{G}(s):=A(s)\,\Gamma(\alpha s+\beta),\qquad\alpha>0,$ (11)
where $A(s)$ is either entire or meromorphic with a pole lattice disjoint from the lattice of $\Gamma(\alpha s+\beta)$.
Observe that a trial state is finite Prony by construction and the Gamma factor of the trial state produces a reference pole lattice
$\Lambda_{G}=\Bigl\{-\frac{\beta+m}{\alpha}\;:\;m=0,1,2,\dots\Bigr\},\qquad\Delta_{G}=\frac{1}{|\alpha|}.$
The argument of the main Theorem can be illustrated in Fig. 1.
Figure 1: Schematic pole-lattice (“ladder”) structure in the Mellin constitutive identity $\tilde{G}(s)=\tilde{K}(s)\,\tilde{H}(-s)$. Crosses mark the simple poles of the kernel factor $\tilde{K}(s)$, located at the integer lattice $s\in\mathbb{Z}$. Filled dots mark a representative arithmetic pole ladder of $\tilde{H}(-s)$ lying on the real axis. Red circles highlight an aligned sublattice where poles of $\tilde{H}(-s)$ coincide with kernel poles; along such an aligned ladder the residue matching yields the first-order recurrence of Theorem 1.
Finite Prony exactness as a relaxation-time question
The trial-state ansatz (11) arises as the Mellin transform (w.r.t. the physical time $t$) of the subordination over relaxation timescales
$G(t)=\int_{0}^{\infty}a(t/\tau)\,\tau^{-(\beta+1)/\alpha}\,\frac{d\tau}{\tau},$
between a finite Prony series $a(t/\tau)$ and the power-law kernel $\tau^{-(\beta+1)/\alpha}$ whose Mellin transform is
$\mathcal{M}\{\tau^{-(\beta+1)/\alpha}\}(s)=\frac{1}{\alpha}\Gamma(\alpha s+\beta).$
Thus the trial state factorization $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ emerges from a multiplicative convolution in the relaxation timescale-space $\tau$. Finite Prony exactness then reduces to the existence of (i) a finite Prony sum $a(t/\tau)$ whose Mellin transform recovers the prescribed modulation $A(s)$ and (ii) the existence of power-law whose subordination yields $G(t)$ on the analyticity strip. This combination amounts to the fulfillment of the two technical conditions identified in Th. 1. Such an interpretation provides the analytic foundation for the ontological reality of finite-mode relaxation dynamics, emerging from the continuous power-law spectrum, through subordination in relaxation timescale space. Mutatis mutandis, if $a(t/\tau)$ is instead allowed to be an infinite Prony series
$a(u)=\sum_{k=1}^{\infty}c_{k}e^{-u/\tau_{k}},$
the resulting representation becomes the infinite Prony ladder constructed in Th. 2, which we call also a transcendental Prony representation.
Theorem 1 (Finite-Prony Representation Criterion).
Let $G^{*}(\omega)$ be a complex viscoelastic modulus whose Mellin transform admits a trial state factorization $\widetilde{G}(s)=A(s)\,\Gamma(\alpha s+\beta)$ as in Definition 2. Denote the pole lattice of the Gamma factor by
$\Lambda_{G}:=\left\{-\frac{\beta+m}{\alpha}:m=0,1,2,\dots\right\}.$
Additionally assume the following genericity conditions:
  1. 1.
    All poles of $\widetilde{G}(s)$ are simple.
  2. 2.
    The only pole coincidences between $\tilde{G}(s)$ and $\tilde{K}(s)\tilde{H}(-s)$ are those forced by the lattice‑alignment condition.
Then, for the constitutive equation
$\tilde{G}(s)=\tilde{K}(s)\,\tilde{H}(-s),\qquad\tilde{K}(s)=-\pi e^{i\pi(1-s)/2}\csc(\pi s),$
a solution $\widetilde{H}(s)\in\mathcal{P}$ exists if and only if:
  1. 1.
    Lattice‑Alignment Condition: For every pole $\lambda$ of $\widetilde{H}(s)$,
[…]
  2. 2.
    Residue‑Compatibility Condition: The residues $\rho_{m}=\operatorname{Res}_{s=-\lambda-\frac{\beta+m}{\alpha}}\widetilde{H}(-s)$ satisfy a decoupled first‑order recurrence
    $\rho_{m+1}=C_{m}\rho_{m},$
    with coefficients $C_{m}$ determined solely by $\tilde{K}(s)$ and $A(s)$, independent of residues from other pole families.
When these conditions hold, the solution $\widetilde{H}(s)\in\mathcal{P}$ is unique up to a scale factor.
Proof.
The proof proceeds in five steps, carefully accounting for the pole structure of both sides of the constitutive equation.
Step 1: Notation and preliminary setup
Observe that $\tilde{K}(s)$ is meromorphic with:
  • 1.
    Simple poles at all integers $s=n\in\mathbb{Z}$ with residues
    $\operatorname{Res}_{s=n}\tilde{K}(s)=\pi ie^{-i\pi n/2}\cdot(-1)^{n}=\pi i(-1)^{n}e^{-i\pi n/2}.$
  • 2.
    An entire factor $e^{-i\pi s/2}$ that does not affect pole locations.
Let $\tilde{H}(s)=H(s)Q(s)$ as in Definition 3. The poles of $\tilde{H}$ come only from $H(s)$. Since $H(s)$ is a ratio of finitely many Gamma functions, its poles lie on a finite union of arithmetic progressions:
$\operatorname{Poles}(H)=\bigcup_{j=1}^{J}\Lambda_{j},\qquad\Lambda_{j}=\Bigl\{\lambda_{j}^{(k)}=-\frac{b_{j}+k}{a_{j}}\;:\;k=0,1,2,\dots\Bigr\}.$
Consequently, $\tilde{H}(-s)$ has poles at
$\operatorname{Poles}(\tilde{H}(-s))=\bigcup_{j=1}^{J}(-\Lambda_{j})=\bigcup_{j=1}^{J}\Bigl\{\frac{b_{j}+k}{a_{j}}\;:\;k=0,1,2,\dots\Bigr\}.$
Let $R_{j,k}:=\operatorname{Res}_{s=\lambda_{j}^{(k)}}\tilde{H}(s)=\operatorname{Res}_{s=\lambda_{j}^{(k)}}H(s)\cdot Q(\lambda_{j}^{(k)})$. Then
$\operatorname{Res}_{s=-\lambda_{j}^{(k)}}\tilde{H}(-s)=-R_{j,k}.$
Step 2: Partial fraction expansion of $K(s)\tilde{H}(-s)$
The product $\tilde{K}(s)\tilde{H}(-s)$ has poles from two sources:
  1. 1.
    At each $\lambda_{j}^{(k)}\in\operatorname{Poles}(\tilde{H}(-s))$, coming from $\tilde{H}(-s)$.
  2. 2.
    At each integer $n\in\mathbb{Z}$, coming from $\tilde{K}(s)$.
Near a pole $s_{0}=-\lambda_{j}^{(k)}$ we have
$\tilde{H}(-s)=\frac{-R_{j,k}}{s-s_{0}}+O(1).$
Multiplying by $K(s)$ gives the principal part
$\frac{-R_{j,k}K(s_{0})}{s-s_{0}}.$
Near an integer pole $n\in\mathbb{Z}$ we have
$\tilde{K}(s)=\frac{\pi i(-1)^{n}e^{-i\pi n/2}}{s-n}+O(1).$
Multiplying by $\tilde{H}(-s)$ (which is regular at $s=n$ unless $n$ coincidentally equals some $\lambda_{j}^{(k)}$) gives the principal part
$\frac{\pi i(-1)^{n}e^{-i\pi n/2}\tilde{H}(-n)}{s-n}.$
For clarity we first treat the noncoincident situation in which $\lambda_{j}^{(k)}\notin\mathbb{Z}$ for all pole locations $\lambda_{j}^{(k)}$ of $\tilde{H}$, so that the poles contributed by $K(s)$ and by $\tilde{H}(-s)$ do not overlap. The coincident-pole cases can be handled separately by local Laurent expansion at the overlap points, and they do not affect the pole-lattice classification discussed below.
Thus the partial fraction expansion is
$K(s)\tilde{H}(-s)=E(s)-\sum_{j=1}^{J}\sum_{k=0}^{\infty}\frac{R_{j,k}K(-\lambda_{j}^{(k)})}{s+\lambda_{j}^{(k)}}+\sum_{n\in\mathbb{Z}}\frac{\pi i(-1)^{n}e^{-i\pi n/2}\tilde{H}(-n)}{s-n},$ (12)
where $E(s)$ is entire.
Step 3: Structure of the left‑hand side $\tilde{G}(s)$
By hypothesis, $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ with $\alpha\neq 0$. The Gamma factor has simple poles at
$s_{m}:=-\frac{\beta+m}{\alpha},\qquad m=0,1,2,\dots,$
with residues
$\operatorname{Res}_{s=s_{m}}\Gamma(\alpha s+\beta)=\frac{(-1)^{m}}{\alpha\,m!}.$
Hence
$\tilde{G}(s)=\sum_{m=0}^{\infty}\frac{A(s_{m})(-1)^{m}/(\alpha\,m!)}{s-s_{m}}+\text{entire part}.$
Let $\rho_{m}:=A(s_{m})(-1)^{m}/(\alpha\,m!)$ denote the residue at $s_{m}$.
Step 4: Necessity of the lattice‑alignment condition
For the identity
$\tilde{G}(s)=\tilde{K}(s)\tilde{H}(-s)$
to hold meromorphically, every pole $s_{m}$ of $\tilde{G}$ must be matched by a pole of $K(s)\tilde{H}(-s)$.
Consider a pole family of $\tilde{H}(s)$ with representative pole $\lambda$. The corresponding points
$s_{n}=-\lambda-\frac{\beta+n}{\alpha},\qquad n=0,1,2,\dots$
are poles of $\tilde{G}(s)$ (from the $\Gamma(\alpha s+\beta)$ factor).
For each such $s_{n}$ to be a pole of the right-hand side $K(s)\tilde{H}(-s)$, the point $s_{n}$ must either:
  • 1.
    Be an integer (so it lies in $\operatorname{Poles}(K)=\mathbb{Z}$), or
  • 2.
[…]
$s_{n}\in\mathbb{Z}\quad\text{or}\quad-s_{n}\in\operatorname{Poles}(\tilde{H})\quad\text{for all }n.$
This establishes the Lattice‑Alignment Condition.
Step 5: Necessity of the residue‑compatibility condition
Assume the Lattice Alignment Condition holds. Focus on a specific aligned sublattice with points $s_{n}=-\lambda-\frac{\beta+n}{\alpha}$.
Equating residues at $s_{n}$ from both sides of the constitutive equation gives:
$\rho_{n}=-R_{j,k_{n}}\tilde{K}(s_{n})+\delta_{n},$ (13)
where $k_{n}$ is such that $s_{n}=-\lambda_{j}^{(k_{n})}$, and $\delta_{n}$ collects contributions from the integer‑pole terms if $s_{n}$ happens to be an integer, plus possible contributions from other pole families if $s_{n}$ coincidentally equals some $\lambda_{j^{\prime}}^{(k^{\prime})}$ with $j^{\prime}\neq j$.
For generic parameters, accidental coincidences beyond those forced by alignment do not occur. Coincidences between distinct arithmetic pole lattices of $\tilde{H}(-s)$ are either finite or occur on a structured subprogression (when lattice spacings are commensurate); thus, apart from finitely many indices (or a periodic subset), the residue matching at $s_{n}$ involves only one pole family.
Moreover, if $s_{n}$ is an integer, the contribution from the integer‑pole series exactly cancels the corresponding part of the $\tilde{H}(-s)$ term, leaving $\delta_{n}=0$. Thus (13) simplifies to
$\rho_{n}=-R_{j,k_{n}}\tilde{K}(s_{n}).$
Now, $K(s_{n})=\frac{\pi ie^{-i\pi s_{n}/2}}{\sin(\pi s_{n})}$ and $\rho_{n}=A(s_{n})(-1)^{n}/(\alpha n!)$. Therefore
$R_{j,k_{n+1}}=\frac{A(s_{n+1})}{A(s_{n})}\cdot\frac{(-1)^{n+1}/(n+1)!}{(-1)^{n}/n!}\cdot\frac{\tilde{K}(s_{n})}{\tilde{K}(s_{n+1})}\cdot R_{j,k_{n}}.$ (14)
Equation (14) is a recurrence along the aligned sublattice. If residues from different pole families couple (i.e., if $\delta_{n}$ depends on $R_{j^{\prime},k^{\prime}}$ for $j^{\prime}\neq j$), then (14) becomes a coupled system involving finitely many parameters $R_{j,k}$ but infinitely many equations. Such a system is infinitely overdetermined and generically has only the trivial solution.
Hence, for a non‑trivial solution to exist, the recurrence must decouple: each aligned sublattice must yield an independent first‑order recurrence involving only residues from that family. This establishes the Residue‑Compatibility Condition.
Step 6: Sufficiency of the conditions
Assume both conditions hold. Then:
  1. 1.
    For each aligned sublattice, (14) gives an explicit first‑order recurrence:
    $R_{j,k_{n+1}}=C_{n}R_{j,k_{n}},\quad C_{n}=\frac{A(s_{n+1})}{A(s_{n})}\cdot\frac{(-1)^{n+1}/(n+1)!}{(-1)^{n}/n!}\cdot\frac{\tilde{K}(s_{n})}{\tilde{K}(s_{n+1})}.$
  2. 2.
    The recurrence can be solved:
    $R_{j,k_{n}}=R_{j,k_{0}}\prod_{\ell=0}^{n-1}C_{\ell}.$
  3. 3.
    The factor $Q(s)$ is determined by requiring that
    $R_{j,k_{n}}=\bigl[\operatorname{Res}_{s=\lambda_{j}^{(k_{n})}}H(s)\bigr]\cdot Q(\lambda_{j}^{(k_{n})})$
    matches the solution. Since $Q(s)$ is of the prescribed entire form, its values on the arithmetic progression $\lambda_{j}^{(k_{n})}$ satisfy a finite‑order recurrence; matching determines the parameters of $Q(s)$.
  4. 4.
    The process involves only algebraic equations for the finite number of parameters in $H(s)$ and $Q(s)$. Thus a solution $\tilde{H}(s)\in\mathcal{P}$ exists and is unique up to an overall scale, which is fixed by the physical balance constraint
    $\int_{0}^{\infty}H(\tau)\frac{d\tau}{t}=G_{0}-G_{\infty}.$
Step 7: Failure implies no solution in $\mathcal{P}$
If the lattice‑alignment condition fails, there exists an infinite set of poles $s_{m}$ of $\tilde{G}$ that are not poles of $K(s)\tilde{H}(-s)$. For an equality to hold, these would need to be cancelled by opposite-sign residues from poles at other locations—an infinite set of linear conditions on finitely many parameters $R_{j,k}$. The only solution is $R_{j,k}=0$ for all $j,k$, giving $\tilde{H}\equiv 0$.
If alignment holds but residue‑compatibility fails, the recurrence equations couple residues from different families, again yielding an infinitely overdetermined system possessing only the trivial solution.
Thus, if either condition fails, no non‑trivial $\tilde{H}\in\mathcal{P}$ satisfies the constitutive equation. ∎
Remark 2.
The phase factor $e^{-i\pi s/2}$ in $K(s)$ affects the coefficients $C_{n}$ in the recurrence but does not alter the pole‑lattice structure. Different branch choices for the logarithm in $(i\omega)^{-s}$ correspond to different representations of the same physical modulus, related by analytic continuation.
5 Transcendental Prony Representation
The classification induced by Th. 1 is completed by the following result.
Theorem 2 (Infinite Prony ladder representation).
Suppose that a given $G(t)$ is completely monotone, but its complex modulus $G^{*}(\omega)$ does not admit a finite Prony representation, that is no $\tilde{H}(s)\in\mathcal{P}$ satisfies the constitutive equation $\tilde{G}(s)=K(s)\tilde{H}(-s)$. Then $G(t)$ admits an exact representation by the infinite Prony ladder:
$H(\tau)=\sum_{k=-\infty}^{\infty}g_{k}\tau_{k}\delta(\tau-\tau_{k}),\qquad\tau_{k}=\tau_{0}q^{k},\;q>1,$
where the weights $g_{k}$ are determined by the continuous relaxation spectrum $H_{c}(\tau)$ existing by Bernstein’s theorem, and the corresponding generalized Mellin kernel satisfies the constitutive equation distributionally. Such representation is called transcendental.
Proof.
By Bernstein’s theorem[11], complete monotonicity of $G(t)$ ensures the existence of a continuous, non‑negative spectrum $H_{c}(\tau)$, such that
$G(t)=G_{\infty}+\int_{0}^{\infty}e^{-t/\tau}H_{c}(\tau)\frac{d\tau}{\tau}.$
Choose a geometric progression $\tau_{k}=\tau_{0}q^{k}$ ($q>1$) and set
$g_{k}=H_{c}(\tau_{k})\ln q.$
Define the discrete approximations
$H_{N}(\tau)=\sum_{k=-N}^{N}g_{k}\tau_{k}\delta(\tau-\tau_{k}).$
For the Mellin transform, standard quadrature theory on a logarithmic grid gives
$\lim_{N\to\infty}\mathcal{M}\{H_{N}\}(s)=\mathcal{M}\{H_{c}\}(s)=\tilde{H}_{c}(s)$
for $s$ in the common strip of convergence. Since $\tilde{G}(s)=K(s)\tilde{H}_{c}(-s)$ holds for the continuous spectrum, and $\mathcal{M}\{H_{N}\}(s)\to\tilde{H}_{c}(s)$ pointwise, the discrete spectra $H_{N}$ converge distributionally to a representation of $H_{c}(\tau)$ as an infinite sum of Dirac masses—the infinite Prony ladder. In the limit, this ladder satisfies the constitutive equation in the sense of distributions. ∎
Corollary 1 (Entire non-exponential spectra are transcendental).
Let $\tilde{H}(s)=\exp(P(s))$ with $P(s)$ be an entire function that is not affine $(P(s)\neq\alpha s+\beta)$. Then the corresponding relaxation spectrum $H(\tau)$ is not finitely Prony‑representable.
In particular, any spectrum whose Mellin transform is an entire function of order greater than 1 (e.g., Gaussian $\exp(\gamma s^{2})$, $\exp(e^{s})$, etc.) requires an infinite Prony ladder for exact representation.
Proof.
First, note that by Proposition 1, a relaxation spectrum is a finite Prony series if and only if its Mellin transform is a finite exponential sum.
Suppose $\exp(P(s))$ were a finite exponential sum: $\exp(P(s))=\sum_{k=1}^{N}c_{k}e^{\lambda_{k}s}$ with distinct $\lambda_{k}$. Taking the logarithmic derivative gives
$P^{\prime}(s)=\frac{\sum_{k}c_{k}\lambda_{k}e^{\lambda_{k}s}}{\sum_{k}c_{k}e^{\lambda_{k}s}}.$
The right-hand side is a rational function of $e^{\lambda_{1}s},\dots,e^{\lambda_{N}s}$. If $P^{\prime}(s)$ is entire, the denominator must have no zeros; otherwise $P^{\prime}(s)$ would have poles. Hence the denominator is a non‑vanishing entire function of $e^{\lambda_{1}s},\dots,e^{\lambda_{N}s}$, which forces it to be constant. Consequently $P^{\prime}(s)$ itself is constant, so $P(s)=Cs+D$. Thus $\exp(P(s))=e^{D}e^{Cs}$ is a single exponential term—a one‑term Prony series.
Therefore, if $P(s)$ is not linear, $\exp(P(s))$ is not a finite exponential sum, and by Proposition 1, $H(\tau)$ cannot be a finite Prony series. By Theorem 1, such spectra fail the residue‑compatibility condition and belong to the transcendental class. Their exact representation requires an infinite Prony ladder, which can be constructed via logarithmic discretization of the continuous spectrum as established by Th. 2. ∎
5.1 Matching of arithmetic progressions
Before stating the classification results, we need a technical result on the intersection of arithmetic progressions.
Lemma 1 (Mismatch of arithmetic progressions).
Let $\alpha>0$, $\beta\in\mathbb{C}$, and define the arithmetic progression
$\Lambda_{\alpha,\beta}=\Bigl\{s_{m}=-\frac{\beta+m}{\alpha}\;:\;m=0,1,2,\dots\Bigr\}.$
Let $\mathbb{Z}$ denote the set of integers. Then:
  1. 1.
    If $\alpha$ is irrational, $\Lambda_{\alpha,\beta}\cap\mathbb{Z}$ contains at most one point.
  2. 2.
    If $\alpha=p/q$ with $p,q\in\mathbb{N}$, $\gcd(p,q)=1$ and $p\neq q$, then $\Lambda_{\alpha,\beta}\cap\mathbb{Z}$ is either empty or an arithmetic progression with spacing p/|pq|p/|p-q|, but in any case $\Lambda_{\alpha,\beta}\not\subseteq\mathbb{Z}$.
  3. 3.
    If $\alpha=1$ (the integer‑spacing case), then $\Lambda_{1,\beta}\subseteq\mathbb{Z}$ if and only if $\beta\in\mathbb{Z}$.
  4. 4.
    In all cases, $\Lambda_{\alpha,\beta}\subseteq\mathbb{Z}$ holds only when $\alpha=1$ and $\beta\in\mathbb{Z}$.
Proof.
For $s_{m}$ to be an integer we need $\beta-m=\alpha k$ for some integer $k$, i.e.
$m+\alpha k=-\beta.$
If $\alpha$ is irrational, this linear Diophantine equation in $m,k$ can have at most one solution.
If $\alpha=p/q$ rational, the equation becomes $qm+pk=-q\beta$. This is a linear Diophantine equation in $m,k$. Its solutions, if any, form an arithmetic progression with step $\operatorname{lcm}(q,p)/q=p/\gcd(p,q)=p$ in $m$. Thus the integers $s_{m}$ that appear are spaced by $p$, not 1, so the progression cannot cover all integers unless $p=1$.
If $\alpha=1$, then $s_{m}=-\beta-m$. These are all integers precisely when $\beta$ is an integer.
The final statement follows from the previous cases: only when $\alpha=1$ can the progression have spacing 1, and only when $\beta\in\mathbb{Z}$ does it lie exactly on the integers. ∎
5.2 Coalescent Poles Cases
When the genericity conditions of Theorem 1 are violated—specifically, when parameters are fine‑tuned so that $\frac{\beta+m}{\alpha}\in\mathbb{Z}$ coincides accidentally with a pole of $\widetilde{H}(-s)$—the product $K(s)\widetilde{H}(-s)$ develops higher‑order poles. The analysis of such coalescent cases follows from Lemma 6:
Proposition 2 (Coalescent poles yield logarithmic kernels).
Suppose $K(s)$ and $\widetilde{H}(-s)$ both have simple poles at $s_{0}$ with residues $a_{-1}$ and $b_{-1}$ respectively. Then:
  1. 1.
    If $a_{-1}b_{-1}\neq 0$, $K(s)\widetilde{H}(-s)$ has a double pole at $s_{0}$ with Laurent expansion
    $K(s)\widetilde{H}(-s)=\frac{a_{-1}b_{-1}}{(s-s_{0})^{2}}+\frac{a_{-1}b_{0}+a_{0}b_{-1}}{s-s_{0}}+O(1).$
  2. 2.
    The inverse Mellin transform of such a double pole produces time‑domain kernels containing logarithmic terms, e.g.,
    $\mathcal{M}^{-1}\left\{\frac{1}{(s-s_{0})^{2}}\right\}(t)\propto t^{-s_{0}}\ln t.$
  3. 3.
    These logarithmic kernels lie outside the finite Prony class $\mathcal{P}$ but admit exact representation as infinite Prony ladders by Th. 2.
Remark 3 (Non-generic coalescent cases).
Coalescent poles—where $\tilde{K}(s)$ integer poles coincide with $\tilde{H}(-s)$ Gamma lattice poles—mark the structurally unstable boundary between finite Prony (rational $G^{*}$) and transcendental (infinite ladder) regimes. Such fine-tuned Diophantine conditions $\frac{\beta+m}{\alpha}=\frac{b_{j}+k}{a_{j}}$ generate logarithmic kernels ($t^{-1}\\ lnt$) requiring infinite Prony series, analyzed via Appendix Lemma 6.
Theorem 3 (Coalescent poles obstruct finite Prony representability).
Let $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ satisfy the hypotheses of Theorem 1 and suppose there exists a coalescent point $s_{0}=-\frac{\beta+m}{\alpha}\in\mathbb{Z}$ for some $m\geq 0$, $s_{0}\neq 0$.
Then, generically (for an open dense set of parameters $\alpha,\beta$ and meromorphic functions $A(s)$), no spectrum $\tilde{H}(s)\in\mathcal{P}$ exists satisfying the constitutive equation
$\tilde{G}(s)=\tilde{K}(s)\,\tilde{H}(-s),\qquad\tilde{K}(s)=-\pi e^{i\pi(1-s)/2}\csc(\pi s).$
Proof.
The proof is constructed by contradiction. Assume, that such $\tilde{H}(s)\in\mathcal{P}$ exists.
1. Functional equation.
Substituting into the constitutive equation gives
$A(s)\Gamma(\alpha s+\beta)=\tilde{K}(s)\,\frac{\prod_{i=1}^{N}\Gamma(-a_{i}s+b_{i})}{\prod_{j=1}^{M}\Gamma(-c_{j}s+d_{j})}\;Q(-s).$ (15)
Divide both sides by $\Gamma(\alpha s+\beta)Q(-s)$ (the division is legitimate outside the poles of these factors) to obtain
$\frac{A(s)}{Q(-s)}=\tilde{K}(s)\,\frac{\prod_{i=1}^{N}\Gamma(-a_{i}s+b_{i})}{\prod_{j=1}^{M}\Gamma(-c_{j}s+d_{j})\,\Gamma(\alpha s+\beta)}.$ (16)
2. Pole structure of the right‑hand side.
The kernel factor $K(s)=-\pi e^{i\pi(1-s)/2}\csc(\pi s)$ possesses simple poles at every integer $s=n\in\mathbb{Z}$. The Gamma factors contribute additional singularities:
  • 1.
    Numerator factors $\Gamma(-a_{i}s+b_{i})$ have poles at $s=(b_{i}+k)/a_{i}$ ($k=0,1,2,\dots$).
  • 2.
    Denominator factors $\Gamma(-c_{j}s+d_{j})$ and $\Gamma(\alpha s+\beta)$ have poles at $s=(d_{j}+k)/c_{j}$ and $s=-(\beta+k)/\alpha$, respectively.
A pole in a denominator Gamma makes the whole ratio in (16) vanish at that point, because $\Gamma(z)$ has no zeros.
3. Cancellation requirement.
The left‑hand side of (16), $\frac{A(s)}{Q(-s)}$, is a meromorphic function. By hypothesis, $A(s)$ has poles disjoint from those of $\Gamma(\alpha s+\beta)$; moreover, $Q(s)=\exp(P(s))$ is entire and non‑vanishing, so Q(s)Q(-s) has no poles and no zeros. Consequently, $\frac{A(s)}{Q(-s)}$ possesses only at most finitely many poles (exactly the poles of $A(s)$ if such exist).
The right‑hand side must therefore also have only finitely many poles. Since $K(s)$ introduces a simple pole at every integer $n\in\mathbb{Z}$, those poles must be cancelled by zeros coming from the denominator Gamma factors.
4. Structural impossibility of the required cancellation.
For the right‑hand side of (16) to have only finitely many poles, the simple poles of $\csc(\pi s)$ in $\tilde{K}(s)$ must be cancelled at all but finitely many integers $n\in\mathbb{Z}$.
Cancellation at $s=n$ occurs when the denominator Gamma factors $\{\Gamma(-c_{j}s+d_{j})\}_{j=1}^{M}$ or $\Gamma(\alpha s+\beta)$ have at least a simple pole there, creating a zero in the ratio of order at least 1. Explicitly, this requires:
$c_{j}n+d_{j}\in\mathbb{Z}_{\leq 0}\quad\text{for some }j,\qquad\text{or}\qquad\alpha n+\beta\in\mathbb{Z}_{\leq 0}.$
Equivalently, $n$ must belong to one of the arithmetic progressions
$\Lambda_{j}:=\Bigl\{\frac{d_{j}+k}{c_{j}}:k=0,1,2,\dots\Bigr\},\qquad\Lambda_{\alpha}:=\Bigl\{-\frac{\beta+k}{\alpha}:k=0,1,2,\dots\Bigr\}.$
The spacing of $\Lambda_{j}$ is $1/c_{j}$, and that of $\Lambda_{\alpha}$ is $1/\alpha$.
The union $\bigl(\bigcup_{j=1}^{M}\Lambda_{j}\bigr)\cup\Lambda_{\alpha}$ must contain all but finitely many integers. A finite union of one-sided arithmetic progressions with fixed spacings can cover all but finitely many integers only if every spacing equals 1 and the progressions together tile $\mathbb{Z}$ exactly. Indeed:
  • 1.
    If any spacing $1/c_{j}$ or $1/\alpha$ exceeds 1, by Lemma 1the corresponding progression cannot be a subset of the integers, and in fact has gaps of size at least 1; consequently, the union of such progressions leaves infinitely many integers uncovered.
  • 2.
    If all spacings equal 1, then each $\Lambda_{j}$ and $\Lambda_{\alpha}$ consists of consecutive integers starting from some initial value. To cover $\mathbb{Z}$ from some point onward, these initial values must form a perfect partition of $\mathbb{Z}$ into intervals.
The progressions $\Lambda_{j},\Lambda_{\alpha}$ are nowhere dense subsets of $\mathbb{R}$ (as subsets of the totally disconnected set $\mathbb{Z}$).
Consequently, for generic choices of $\alpha,\beta,\{c_{j},d_{j}\}$–and in particular when $\alpha\neq 1$—the required cancellation fails at infinitely many integers. The right‑hand side of (16) then possesses infinitely many poles, contradicting the left‑hand side’s finite poles.
5. Even if cancellation were arranged, residue matching fails.
Suppose, non‑generically, that the denominator Gamma factors indeed produce zeros at every unwanted integer. At the surviving poles (those in $\Lambda_{G}$) the residues of both sides of (15) must coincide. This yields an infinite set of equations for the finite collection of parameters $\{a_{i},b_{i},c_{j},d_{j}\}$ and the entire function $P(s)$ defining $Q(s)$. For generic $A(s),\alpha,\beta$ this infinitely overdetermined system has no solution.
Since the necessary cancellations and residue compatibilities cannot be satisfied simultaneously for generic parameters, no $\widetilde{H}(s)\in\mathcal{P}$ can satisfy (15). Hence the constitutive equation admits no solution within the finite Prony class $\mathcal{P}$. ∎
6 Classification of Common Viscoelastic Models
Tables 1 and 2 summarize the Mellin‑space analysis of standard viscoelastic models. The Lattice‑Alignment Condition of Theorem 1 together with Lemma 1 determines whether a finite Prony representation exists, while the Residue‑Compatibility Condition governs cases with integer spacing $\Delta_{G}=1$.
Table 1: Mellin‑space classification of viscoelastic models: Finite Prony class $\mathcal{P}$.
Model Complex Modulus $G^{*}(\omega)$ $\Delta_{G}$} In $\mathcal{P}$? Reason
Maxwell $G_{\infty}+g\dfrac{i\omega\tau}{1+i\omega\tau}$ 1 Yes Integer lattice; residues decouple
Standard Linear Solid (SLS) $G_{\infty}+\dfrac{g}{1+i\omega\tau}$ 1 Yes Integer lattice; residues decouple
$\Delta_{G}$: spacing of the dominant Gamma‑factor pole lattice in $\tilde{G}(s)$. Finite Prony representability requires $\Delta_{G}=1$ (Lattice‑Alignment) and decoupled residue recurrences (Residue‑Compatibility). Notation: $G_{\infty}$ = high‑frequency (glassy) modulus, $g$ = spring constant, $\tau$ = relaxation time.
Table 2: Mellin‑space classification of viscoelastic models: Transcendental class (infinite Prony ladders).
Model Complex Modulus $G^{*}(\omega)$ $\Delta_{G}$} In $\mathcal{P}$? Obstruction
Power‑law $G_{0}(i\omega\tau_{0})^{\beta},\;0<\beta<1$ No Continuous spectrum (no discrete poles)
Cole–Cole $G_{\infty}+\dfrac{\Delta G}{1+(i\omega\tau)^{-\alpha}}$ $1/\alpha$ No $\Delta_{G}\neq 1$ for $\alpha\neq 1$ (Lemma 1)
Cole–Davidson $G_{\infty}+\dfrac{\Delta G}{(1+i\omega\tau)^{\beta}}$ 1 No Residue‑compatibility fails ($\Gamma(\beta-s)$ coupling)
Havriliak–Negami $G_{\infty}+\dfrac{\Delta G}{[1+(i\omega\tau)^{\alpha}]^{\beta}}$ $\alpha$ No $\Delta_{G}\neq 1$ for $\alpha\neq 1$
Fractional Zener $G_{e}\dfrac{1+(i\omega\tau)^{\alpha}}{1+\delta(i\omega\tau)^{\alpha}}$ $\alpha$ No $\Delta_{G}\neq 1$ for $\alpha\neq 1$
Gaussian $G^{*}(\omega)$ via Gaussian $H(\tau)$ 1 No Entire Mellin symbol; residue condition fails
$\Delta_{G}$: spacing of the dominant Gamma‑factor pole lattice. Transcendental models require infinite Prony ladders (Theorem 2). Notation: $0<\alpha,\beta<1$, $\Delta G=G_{0}-G_{\infty}$, $G_{e}$ = equilibrium modulus, $\delta=G_{e}/G_{\infty}<1$. The Cole–Cole entry uses the viscoelastic sign convention.
Proposition 3.
$G^{*}(\omega)=G_{0}(i\omega\tau_{0})^{\beta}$ ($0<\beta<1$), the Cole–Cole modulus admit no exact representation by a relaxation spectrum in the finite Prony class $\mathcal{P}$, except in the degenerate limits $\beta\to 0,1$ or $\alpha\to 1$.
Proof.
$\tilde{G}(s)=G_{0}\tau_{0}^{\beta}\delta(s-\beta)$ is a Dirac delta, which cannot be expressed as $K(s)\tilde{H}(-s)$ with $\tilde{H}\in\mathcal{P}$ because the right-hand side has poles on integer lattices (from $K(s)$) and possibly other arithmetic progressions, not a single isolated pole.
Proposition 4 (Power‑law, Cole–Cole, Havriliak-Negami and Fractional Zener models are not in $\mathcal{P}$).
Cole–Cole, Havriliak-Negami and Fractional Zener models are not in $\mathcal{P}$ due to Lattice Incompatibility.
Proof.
Cole–Cole: $\tilde{G}(s)\propto\Gamma(s)\Gamma(1-s)/\Gamma(1-\alpha s)$. The poles come from $\Gamma(s)$ at s=ns=-n (spacing 1) and from the zeros of $\Gamma(1-\alpha s)$ at $s=(1+m)/\alpha$ (spacing $1/\alpha$). By Lemma 1, if $\alpha\neq 1$, this progression is not contained in the integers, nor in any finite union of integer-spaced progressions. Hence the Lattice-Alignment Condition fails.
Havriliak–Negami: For the Havriliak–Negami modulus $G^{*}(\omega)=G_{\infty}+\Delta G/[1+(i\omega\tau)^{\alpha}]^{\beta}$, the Mellin transform contains factors $\Gamma(s/\alpha)$ and $\Gamma(\beta-s/\alpha)$. The factor $\Gamma(s/\alpha)$ generates poles at $s=-\alpha m$ ($m=0,1,2,\dots$) with spacing $\alpha$. By Lemma 1, if $\alpha\neq 1$, this arithmetic progression cannot be contained in the integers. Consequently, the Lattice-Alignment Condition of Theorem 1 fails, and no finite Prony representation exists.
Fractional–Zener: For the standard fractional Zener model $G^{*}(\omega)=G_{g}\frac{1+(i\omega\tau)^{\alpha}}{1+\delta(i\omega\tau)^{\alpha}}$, the viscoelastic part can be written as $\Delta G^{*}(\omega)=\Delta G/[1+\frac{1}{\delta}(i\omega\tau)^{-\alpha}]$. Its Mellin transform contains Gamma factors with poles at $s=(1+m)/\alpha$ (spacing $1/\alpha$). By Lemma 1 with the same argument as for Cole–Cole, when $\alpha\neq 1$ the progression with spacing $1/\alpha$ cannot align with the integer lattice. Therefore the Lattice-Alignment Condition fails, precluding finite Prony representability.
In all three models, the failure occurs for any non‑integer $\alpha$; the degenerate cases $\alpha=1$ reduce to classical (finite Prony) models. ∎
7 Test for Finite Prony Representability
Based on Theorem 1, we formulate a step‑by‑step test to determine whether a given complex modulus $G^{*}(\omega)$ admits a finite Prony representation.
  1. 1.
    Compute the Mellin transform. Determine $\tilde{G}(s)=\mathcal{M}\{G^{*}(\omega)-G_{\infty}\}(s)$. Identify any non‑degenerate Gamma factors $\Gamma(\alpha s+\beta)$ with $\alpha\neq 0$. If none exist, the modulus may be rational (finite Prony) or distributional (power‑law).
  2. 2.
    Determine the pole‑lattice spacing. From each Gamma factor, extract the fundamental spacing $\Delta_{G}=1/|\alpha|$.
  3. 3.
    Postulate a candidate spectrum template. Construct a physically motivated $\tilde{H}_{\text{cand}}(s)\in\mathcal{P}$. For a discrete spectrum, this typically involves factors like $\Gamma(s+1)$ (for a simple pole at s=1s=-1) or ratios of Gamma functions with integer parameters.
  4. 4.
    Extract the candidate’s lattice spacings. Analyze the Gamma‑ratio factor of $\tilde{H}_{\text{cand}}(s)$ to obtain its distinct spacings $\Delta_{1},\dots,\Delta_{K}$.
  5. 5.
    Test the necessary alignment condition. A necessary condition for the Lattice‑Alignment Condition to be satisfiable is that $\Delta_{G}$ be an integer combination of the candidate’s spacings:
    $\Delta_{G}=\sum_{i=1}^{K}n_{i}\Delta_{i},\qquad n_{i}\in\mathbb{Z}_{\geq 0}.$
    If this diophantine condition fails, finite representability is impossible.
  6. 6.
    Interpretation
    • (a)
      If the condition fails, the modulus is provably not finitely representable in $\mathcal{P}$ (and hence not by any finite Prony series). It is transcendentally representable, and an infinite ladder construction should be pursued.
    • (b)
      If the condition holds, finite representability is possible but not guaranteed; one must additionally check the Residue‑Compatibility Condition by examining the recurrence along the aligned sublattice.
This test provides a rigorous filter. For example, applying it to the Cole–Cole model ($\Delta_{G}=1/\alpha$) with a candidate having integer spacings $\Delta_{i}=1$ immediately shows that alignment requires $1/\alpha$ to be an integer—impossible for non‑integer $\alpha$. Hence the Cole–Cole model is not finitely representable.
Proposition 5 (Cole–Davidson model is not in $\mathcal{P}$).
The Cole–Davidson modulus $G^{*}(\omega)=G_{\infty}+\Delta G/(1+i\omega\tau)^{\beta}$ with $0<\beta<1$ does not belong to the finite Prony class $\mathcal{P}$, except in the trivial limit $\beta\to 1$ (which reduces to the standard linear solid).
Proof.
The Mellin transform of the Cole–Davidson viscoelastic part is
$\widetilde{G}(s)=\Delta G\,\tau^{-s}i^{-s}\frac{\Gamma(s)\Gamma(\beta-s)}{\Gamma(\beta)},\qquad 0<\Re s<\beta.$
This factors as $\widetilde{G}(s)=A(s)\,\Gamma(s)$ with $A(s)=\Delta G\,\tau^{-s}i^{-s}\Gamma(\beta-s)/\Gamma(\beta)$. Hence in the trial‑state notation $\widetilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ we have $\alpha=1$, $\beta_{\text{trial}}=0$; the pole spacing is $\Delta_{G}=1$.
Although the lattice spacing equals $1$ and therefore satisfies the necessary Lattice‑Alignment Condition, the Residue‑Compatibility Condition of Theorem 1 fails. Indeed, the factor $\Gamma(\beta-s)$ in $A(s)$ introduces poles at $s=\beta+m$ ($m=0,1,2,\dots$) whose residues couple with those coming from $\Gamma(s)$. This coupling yields an infinitely overdetermined system for the residues of any candidate $\widetilde{H}(s)\in\mathcal{P}$; the only solution is the trivial one $\widetilde{H}\equiv 0$.
Consequently, no spectrum $\widetilde{H}(s)$ in the finite Prony class $\mathcal{P}$ can satisfy the constitutive equation $\widetilde{G}(s)=\widetilde{K}(s)\widetilde{H}(-s)$. The Cole–Davidson relaxation spectrum therefore admits only a transcendental (infinite‑ladder) Prony representation, as guaranteed by Theorem 2. ∎
8 Conclusions
In conclusion, we have established a universal Mellin‑space criterion for finite Prony representability of linear viscoelastic moduli. The criterion—comprising the fulfillment of both the Lattice- Alignment Condition (pole‑spacing constraint) and the Residue - Compatibility Condition (decoupled recurrence requirement)—provides a complete analytic characterization of when an analytical viscoelastic modulus admits an exact representation by finitely many relaxation modes.
This framework explains the fundamental dichotomy between classical and fractional viscoelasticity:
  • 1.
    Classical spring‑dashpot networks (Maxwell, SLS) yield rational transfer functions whose Mellin symbols possess integer‑spaced pole lattices, satisfying both conditions and hence belonging to the finite Prony class $\mathcal{P}$.
  • 2.
    Fractional constitutive laws (Cole–Cole, Havriliak–Negami, fractional Zener) produce non‑integer lattice spacings ($\Delta_{G}=\alpha$ or $1/\alpha$ with $\alpha\neq 1$), violating the Lattice‑Alignment Condition by Lemma 1.
  • 3.
    Borderline cases like Cole–Davidson, while possessing integer spacing ($\Delta_{G}=1$), fail the Residue‑Compatibility Condition due to coupling between residue recurrences from distinct Gamma factors.
By extending the analysis to the maximal finite Prony class $\mathcal{P}$, we demonstrate that the obstruction is structural, not merely an artifact of a restricted ansatz. For transcendentally representable moduli, we have provided a constructive logarithmic‑discretization method that yields exact infinite Prony ladders, with the static modulus gap imposing a sum‑rule analogous to a quantization condition.
The Mellin‑space framework unifies the treatment of power‑law, Gaussian, Cole–Cole, Havriliak–Negami, Cole–Davidson, and Mittag‑Leffler models, revealing a direct correspondence between fractional‑order operators and the arithmetic geometry of their kernel singularities.
The present work thereby draws a sharp line between “finitely representable” viscoelastic materials (those in $\mathcal{P}$) and “transcendentally representable” ones (those requiring infinite Prony ladders), offering a new paradigm for classifying, modeling, and decoding linear viscoelastic behavior from complex‑analytic principles.
Acknowledgment
This work was supported by the European Union’s Horizon Europe programme under grant agreement No. 101086815 (VIBTaTE).
Appendix A Mellin Transform of Constitutive Identity
A.1 Derivation of the Mellin-space constitutive equation
Lemma 2 (Mellin transform of the viscoelastic kernel).
For $\tau>0$ and $0<\Re s<1$, the Mellin transform (in $\omega$) of $\frac{i\omega\tau}{1+i\omega\tau}$ is given by the meromorphic continuation
$\mathcal{M}_{\omega}\!\left\{\frac{i\omega\tau}{1+i\omega\tau}\right\}(s)=-\,\tau^{-s}\,i^{-s}\,\frac{\pi}{\sin(\pi s)}.$ (17)
Proof.
With $x=\omega\tau$,
$\int_{0}^{\infty}\frac{i\omega\tau}{1+i\omega\tau}\,\omega^{s-1}\,d\omega=\tau^{-s}\int_{0}^{\infty}\frac{i\,x^{s}}{1+ix}\,dx.$
Let $u=s+1$. By the standard Beta integral (analytic continuation),
$\int_{0}^{\infty}\frac{x^{u-1}}{1+ix}\,dx=i^{-u}\frac{\pi}{\sin(\pi u)}.$
Hence
$\int_{0}^{\infty}\frac{i\,x^{s}}{1+ix}\,dx=i\int_{0}^{\infty}\frac{x^{u-1}}{1+ix}\,dx\Big|_{u=s+1}=-\,i^{-s}\frac{\pi}{\sin(\pi s)},$
which gives (17). ∎
Choose a strip $a<\Re s<b$ with $0<a<b<1$ where the Mellin transforms of $G_{v}$ and $H$ converge absolutely, so that Fubini’s theorem applies. Taking the Mellin transform of the frequency-domain constitutive relation and using Lemma 2 yields
$\tilde{G}(s)=-\,\frac{\pi\,i^{-s}}{\sin(\pi s)}\int_{0}^{\infty}H(\tau)\,\tau^{-s-1}\,d\tau=-\,\frac{\pi\,i^{-s}}{\sin(\pi s)}\,\tilde{H}(-s),\qquad a<\Re s<b.$ (18)
Defining the meromorphic kernel factor
$K(s):=-\,\frac{\pi\,i^{-s}}{\sin(\pi s)}=-\,\frac{\pi e^{-i\pi s/2}}{\sin(\pi s)},$
we obtain the Mellin-space constitutive equation (10).
Appendix B Auxiliary Lemmas for the Proof of Theorem 1
Lemma 3 (Locality of principal parts).
Let $F$ and $G$ be meromorphic in a neighborhood of $s_{0}\in\mathbb{C}$. If $F(s)=G(s)$ for all $s$ in some punctured neighborhood $0<|s-s_{0}|<r$, then the principal parts of $F$ and $G$ at $s_{0}$ coincide. In particular, $F$ has a pole at $s_{0}$ if and only if $G$ has a pole at $s_{0}$, and if the pole is simple then $\operatorname{Res}_{s=s_{0}}F=\operatorname{Res}_{s=s_{0}}G$.
Proof.
Define H(s):=F(s)G(s)H(s):=F(s)-G(s). Then $H$ is meromorphic near $s_{0}$ and satisfies $H(s)=0$ for all $s$ with $0<|s-s_{0}|<r$. If $H$ had a pole at $s_{0}$, it would be unbounded arbitrarily close to $s_{0}$, contradicting $H\equiv 0$ on the punctured neighborhood. Hence $H$ is holomorphic at $s_{0}$. By the identity theorem for holomorphic functions, $H\equiv 0$ on $|s-s_{0}|<r$. Therefore $F$ and $G$ have identical Laurent expansions at $s_{0}$, so their principal parts coincide. ∎
Lemma 4 (Pole containment for the Mellin constitutive product).
Let $K$ and $\tilde{H}$ be meromorphic functions, and define $F(s):=K(s)\tilde{H}(-s)$. Then
[…]
In particular, for $K(s)=\pi e^{i\pi(1-s)/2}/\sin(\pi s)$ one has
[…]
Proof.
[…]
Lemma 5 (Residues away from coincident poles).
Let $K$ be meromorphic and $\tilde{H}$ meromorphic.
  1. 1.
    If $s_{0}\notin\operatorname{Poles}(K)$ $s_{0}\in-\operatorname{Poles}(\tilde{H}), and$ $\tilde{H}(-s)$ has a simple pole at $s_{0}$, then
    $\operatorname{Res}_{s=s_{0}}\bigl(K(s)\tilde{H}(-s)\bigr)=K(s_{0})\,\operatorname{Res}_{s=s_{0}}\tilde{H}(-s).$
  2. 2.
    If $s_{0}\in\operatorname{Poles}(K)$ $s_{0}\notin-\operatorname{Poles}(\tilde{H}), and$ $K$ has a simple pole at $s_{0}$, then
    $\operatorname{Res}_{s=s_{0}}\bigl(K(s)\tilde{H}(-s)\bigr)=\bigl(\operatorname{Res}_{s=s_{0}}K(s)\bigr)\,\tilde{H}(-s_{0}).$
Proof.
(1) Since $K$ is holomorphic at $s_{0}$, write $K(s)=K(s_{0})+O(s-s_{0})$. Since $\tilde{H}(-s)$ has a simple pole at $s_{0}$, write $\tilde{H}(-s)=\dfrac{b_{-1}}{s-s_{0}}+b_{0}+O(s-s_{0})$. Then
$K(s)\tilde{H}(-s)=\frac{K(s_{0})b_{-1}}{s-s_{0}}+O(1),$
so the residue equals $K(s_{0})b_{-1}=K(s_{0})\operatorname{Res}_{s=s_{0}}\tilde{H}(-s)$.
(2) Since $\tilde{H}(-s)$ is holomorphic at $s_{0}$, write $\tilde{H}(-s)=\tilde{H}(-s_{0})+O(s-s_{0})$. Since $K$ has a simple pole at $s_{0}$, write $K(s)=\dfrac{a_{-1}}{s-s_{0}}+a_{0}+O(s-s_{0})$. Then
$K(s)\tilde{H}(-s)=\frac{a_{-1}\tilde{H}(-s_{0})}{s-s_{0}}+O(1),$
so the residue equals $a_{-1}\tilde{H}(-s_{0})=(\operatorname{Res}_{s=s_{0}}K(s))\,\tilde{H}(-s_{0})$. ∎
Lemma 6 (Coincident simple poles produce a double pole unless constrained).
Assume $K$ and $\tilde{H}(-s)$ have simple poles at the same point $s_{0}$. Write the local expansions
$K(s)=\frac{a_{-1}}{s-s_{0}}+a_{0}+O(s-s_{0}),\qquad\tilde{H}(-s)=\frac{b_{-1}}{s-s_{0}}+b_{0}+O(s-s_{0}).$
Then
$K(s)\tilde{H}(-s)=\frac{a_{-1}b_{-1}}{(s-s_{0})^{2}}+\frac{a_{-1}b_{0}+a_{0}b_{-1}}{s-s_{0}}+O(1).$
In particular, $K(s)\tilde{H}(-s)$ has at most a simple pole at $s_{0}$ if and only if $a_{-1}b_{-1}=0$.
Proof.
Multiply the two Laurent expansions:
$\left(\frac{a_{-1}}{s-s_{0}}+a_{0}+O(s-s_{0})\right)\left(\frac{b_{-1}}{s-s_{0}}+b_{0}+O(s-s_{0})\right).$
The coefficient of $(s-s_{0})^{-2}$ is $a_{-1}b_{-1}$, and the coefficient of $(s-s_{0})^{-1}$ is $a_{-1}b_{0}+a_{0}b_{-1}$, which yields the stated principal part. The pole order is $\leq 1$ if and only if the $(s-s_{0})^{-2}$ coefficient vanishes, i.e. $a_{-1}b_{-1}=0$. ∎
Lemma 7 (Intersection of Gamma pole lattices).
Let $a,a^{\prime}>0$ and $b,b^{\prime}\in\mathbb{C}$. Define the (one-sided) arithmetic lattices
$\Lambda(a,b):=\left\{-\frac{b+k}{a}\;:\;k\in\mathbb{Z}_{\geq 0}\right\},\qquad\Lambda(a^{\prime},b^{\prime}):=\left\{-\frac{b^{\prime}+k^{\prime}}{a^{\prime}}\;:\;k^{\prime}\in\mathbb{Z}_{\geq 0}\right\}.$
Then:
  1. 1.
    If $a/a^{\prime}\notin\mathbb{Q}$, then $\Lambda(a,b)\cap\Lambda(a^{\prime},b^{\prime})$ contains at most one point.
  2. 2.
    If $a/a^{\prime}\in\mathbb{Q}$, write $a/a^{\prime}=p/q$ with $p,q\in\mathbb{N}$ and $\gcd(p,q)=1$. Then either $\Lambda(a,b)\cap\Lambda(a^{\prime},b^{\prime})=\varnothing$, or else the set of solutions $(k,k^{\prime})\in\mathbb{Z}_{\geq 0}^{2}$ to
    $\frac{b+k}{a}=-\frac{b^{\prime}+k^{\prime}}{a^{\prime}}$
    is of the form $(k,k^{\prime})=(k_{0},k_{0}^{\prime})+t(p,q)$ for integers $t$ in some one-sided range, and consequently $\Lambda(a,b)\cap\Lambda(a^{\prime},b^{\prime})$ is a (one-sided) arithmetic progression.
Proof.
A point $s$ lies in the intersection if and only if there exist $k,k^{\prime}\in\mathbb{Z}_{\geq 0}$ such that
$\frac{b+k}{a}=-\frac{b^{\prime}+k^{\prime}}{a^{\prime}}.$
Equivalently,
$a^{\prime}(b+k)=a(b^{\prime}+k^{\prime})\quad\Longleftrightarrow\quad a^{\prime}k-ak^{\prime}=ab^{\prime}-a^{\prime}b.$ (19)
(1) Suppose $a/a^{\prime}\notin\mathbb{Q}$ and assume there are two distinct solutions $(k_{1},k_{1}^{\prime})\neq(k_{2},k_{2}^{\prime})$ of (19). Subtracting the two equalities gives
$a^{\prime}(k_{1}-k_{2})=a(k_{1}^{\prime}-k_{2}^{\prime}).$
If $k_{1}\neq k_{2}$ then $a/a^{\prime}=(k_{1}-k_{2})/(k_{1}^{\prime}-k_{2}^{\prime})\in\mathbb{Q}$, a contradiction. If $k_{1}=k_{2}$ then necessarily $k_{1}^{\prime}=k_{2}^{\prime}$, again a contradiction. Hence at most one solution exists, so the intersection contains at most one point.
(2) Now assume $a/a^{\prime}=p/q$ in lowest terms, so $qa=pa^{\prime}$. Multiply (19) by $q$ and substitute $qa=pa^{\prime}$ to obtain
$(qa^{\prime})k-(pa^{\prime})k^{\prime}=q(ab^{\prime}-a^{\prime}b).$
Dividing by $a^{\prime}\neq 0$ gives
$qk-pk^{\prime}=q\left(\frac{a}{a^{\prime}}b^{\prime}-b\right).$ (20)
If (20) has no nonnegative integer solutions, the intersection is empty. If it has a solution $(k_{0},k_{0}^{\prime})$, then any other solution $(k,k^{\prime})$ must satisfy the homogeneous equation $q(k-k_{0})=p(k^{\prime}-k_{0}^{\prime})$. Since $\gcd(p,q)=1$, this implies $k-k_{0}=pt$ and $k^{\prime}-k_{0}^{\prime}=qt$ for some $t\in\mathbb{Z}$. Restricting to $k,k^{\prime}\geq 0$ restricts $t$ to a one-sided integer range. Thus the solutions form $(k,k^{\prime})=(k_{0},k_{0}^{\prime})+t(p,q)$ and the intersection points s=(b+k)/as=-(b+k)/a form a one-sided arithmetic progression. ∎
Lemma 8 (Decoupling necessity for viscoelastic constitutive equation).
Let $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ satisfy the hypotheses of Theorem 1. Suppose that for some pole family $\Lambda_{j}$ of $\tilde{H}(s)$, the residue recurrence along an aligned sublattice couples residues from two distinct pole families $\Lambda_{j}$ and $\Lambda_{k}$ ($k\neq j$) of $\tilde{H}(s)$.
Then for generic parameters $\alpha,\beta$ and generic meromorphic $A(s)$, the only solution of the coupled system is $R_{j}=R_{k}=0$, implying $\tilde{H}\equiv 0$.
Proof.
Consider the Mellin-space constitutive equation $\tilde{G}(s)=\tilde{K}(s)\tilde{H}(-s)$. For an aligned sublattice where poles of $\tilde{H}(-s)$ coincide with poles of $\tilde{K}(s)\tilde{H}(-s)$, the residue matching gives (from the proof of Theorem 1, equation (13)):
$\rho_{n}=-R_{j,k_{n}}\tilde{K}(s_{n})+\delta_{n},$
where $\rho_{n}=\operatorname{Res}_{s=s_{n}}\tilde{G}(s)$, $R_{j,k_{n}}=\operatorname{Res}_{s=s_{n}}\tilde{H}(-s)$ (from the $j$-th family), and $\delta_{n}$ contains contributions from other families if $s_{n}$ coincides with poles from $\Lambda_{k}$.
Suppose $\delta_{n}$ involves residues from family $\Lambda_{k}$: $\delta_{n}=cR_{k,\ell_{n}}$ for some constant $c$ and appropriate index $\ell_{n}$. Then (B.1) becomes:
$\rho_{n}=-R_{j,k_{n}}\tilde{K}(s_{n})+cR_{k,\ell_{n}},\qquad n\in\mathcal{N},$
where $\mathcal{N}$ is an infinite set of indices along the aligned sublattice.
Asymptotic analysis. The known quantities have explicit asymptotics:
  • 1.
    $\rho_{n}=A(s_{n})\frac{(-1)^{n}}{\alpha\,n!}\sim\frac{C(-1)^{n}}{n!}$ (factorial decay)
  • 2.
    $\tilde{K}(s_{n})=-\pi e^{i\pi(1-s_{n})/2}\csc(\pi s_{n})\sim C^{\prime}e^{i\pi s_{n}/2}$ (oscillatory, bounded)
  • 3.
    $R_{j,k_{n}}=[\operatorname{Res}_{s=s_{n}}H(-s)]Q(-s_{n})$ where $H$ is a Gamma-ratio, giving $R_{j,k_{n}}\sim\frac{C_{j}(-1)^{k_{n}}}{k_{n}!}$ (factorial decay)
  • 4.
    Similarly, $R_{k,\ell_{n}}\sim\frac{C_{k}(-1)^{\ell_{n}}}{\ell_{n}!}$ (factorial decay)
The key is that the decay rates of $R_{j,k_{n}}$ and $R_{k,\ell_{n}}$ are both factorial but with potentially different scales. Rearranging (B.2):
$R_{j,k_{n}}\tilde{K}(s_{n})-cR_{k,\ell_{n}}=-\rho_{n}.$
For large $n$, the left side is a linear combination of two factorial-decaying sequences, while the right side is another factorial-decaying sequence with a specific coefficient $A(s_{n})$.
Overdetermination argument. Consider the ratios. Let $r_{n}=R_{j,k_{n}}/R_{k,\ell_{n}}$. Using Gamma-residue formulas:
$r_{n}\sim\frac{C_{j}}{C_{k}}\frac{\Gamma(\ell_{n}+1)}{\Gamma(k_{n}+1)}\sim\frac{C_{j}}{C_{k}}\frac{\ell_{n}!}{k_{n}!}.$
Since $k_{n}$ and $\ell_{n}$ are linearly related (both along arithmetic progressions), say $k_{n}=an+b$, $\ell_{n}=cn+d$, we have by Stirling:
$r_{n}\sim Cn^{(c-a)n}e^{(a-c)n}\quad\text{(super-exponential growth or decay)}.$
Thus either $r_{n}\to 0$ or $r_{n}\to\infty$ super-exponentially. In either case, equation (B.3) divided by the larger of the two residues yields for large $n$:
$\text{(dominant term)}\sim\text{(negligible term)}=-\rho_{n}/\text{(normalization)}.$
But $\rho_{n}$ decays with yet another factorial rate (determined by $\Gamma(\alpha s_{n}+\beta)$). For generic parameters, the three factorial sequences $(\rho_{n},R_{j,k_{n}},R_{k,\ell_{n}})$ have incommensurate decay/growth ratios.
Formally, take three consecutive equations from (B.3) for $n,n+1,n+2$:
$\begin{pmatrix}\tilde{K}(s_{n})&-c&-\rho_{n}\\ \tilde{K}(s_{n+1})&-c&-\rho_{n+1}\\ \tilde{K}(s_{n+2})&-c&-\rho_{n+2}\end{pmatrix}\begin{pmatrix}R_{j,k_{n}}\\ R_{k,\ell_{n}}\\ 1\end{pmatrix}=0.$
For large $n$, the matrix is dominated by the factorial decay in the third column. More precisely, dividing each row by $\rho_{n}$, $\rho_{n+1}$, $\rho_{n+2}$ respectively yields a system where the coefficients of $R_{j,k_{n}}$ and $R_{k,\ell_{n}}$ become super-exponentially large or small relative to 1. Such a system can only be satisfied if $R_{j,k_{n}}=R_{k,\ell_{n}}=0$ for all large $n$, implying all residues vanish.
Thus the only solution to the coupled recurrence is the trivial one. For a non‑trivial $\tilde{H}(s)$ to exist, the recurrence must decouple: each aligned sublattice must involve residues from only one pole family. ∎
Appendix C Auxiliary Lemmas for the Proof of Theorem 3
Lemma 9 (Necessary condition for coalescent poles).
Let $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ with a coalescent point $s_{0}=-\frac{\beta+m}{\alpha}\in\mathbb{Z}$ for some $m\geq 0$.
If there exists $\tilde{H}(s)\in\mathcal{P}$ satisfying the constitutive equation $\tilde{G}(s)=\tilde{K}(s)\tilde{H}(-s)$, then:
  1. 1.
    $\widetilde{H}(-s)$ must be analytic at $s=s_{0}$.
  2. 2.
    The residues satisfy:
    $\operatorname{Res}_{s=s_{0}}\tilde{G}(s)=(\operatorname{Res}_{s=s_{0}}\tilde{K}(s))\cdot\tilde{H}(-s_{0}).$
Proof.
At $s=s_{0}$:
  • 1.
    $\tilde{G}(s)$ has a simple pole (from $\Gamma(\alpha s+\beta)$)
  • 2.
    $\tilde{K}(s)$ has a simple pole (since all integers are poles of $K$)
Suppose $\tilde{H}(-s)$ also has a pole at $s_{0}$. Then by Lemma 6, $\tilde{K}(s)\tilde{H}(-s)$ has at least a double pole at $s_{0}$, while $\widetilde{G}(s)$ has only a simple pole. This contradicts the equality.
Therefore, $\tilde{H}(-s)$ must be analytic at $s_{0}$. Writing Laurent expansions:
$\tilde{G}(s)=\frac{\rho_{G}}{s-s_{0}}+O(1),\quad\tilde{K}(s)=\frac{\rho_{K}}{s-s_{0}}+O(1),\quad\tilde{H}(-s)=h_{0}+O(s-s_{0}),$
we have:
$\tilde{K}(s)\tilde{H}(-s)=\frac{\rho_{K}h_{0}}{s-s_{0}}+O(1).$
Equating with $\tilde{G}(s)$ gives $\rho_{G}=\rho_{K}h_{0}$. ∎